OpenStudy (anonymous):

If the radius of a circle is multiplied by 4, then how many times larger is the circumference of the new circle?

OpenStudy (anonymous):

The circumference of a circle is given by \[2*\pi*r\]

OpenStudy (anonymous):

and lets say we pick r=1 we get \[2*\pi\]

OpenStudy (anonymous):

now lets say that we multiply the radius by 4 so now r=4

OpenStudy (anonymous):

\[8*\pi\]

OpenStudy (anonymous):

8 pi is 4 times bigger than 2 pi

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

Look at my last post

OpenStudy (anonymous):

so what do i do next

OpenStudy (anonymous):

so is 32

OpenStudy (anonymous):

wait can you tell me what the original radius equals?

OpenStudy (anonymous):

i really dont know

OpenStudy (anonymous):

Answer is 4

OpenStudy (anonymous):

thanks